A local government operates a small park near to its city hall. The department uses its attendance to measure its workload in determining the budget for the park. The city manager has long believed that the park has too few visitors and the department should reach out to more customers or its budget will need to be reduced. You are an analyst for the city. The city manager has asked you to prepare an analysis for the park to determine whether the budget request for the park is justified in the city’s budget proposal. You randomly selected 25 days in the past year and calculated the park attendance data: 5, 3, 10, 1, 2, 3, 4, 3, 5, 100, 4, 3, 2, 4, 25, 150, 3, 3, 5, 4, 8, 7, 10, 15, and 30.

Write a paragraph to explain the meaning of these statistics. (10 points)

Do you recommend the use of the mean in your presentation? Why or why not? ( 10 points)

Do you recommend the use of the median or the mode? Why or why not? (10 points)
Answer #1

The data collected is about the attendance of people who come to park in past years for 25 random days.

The statistic we can use here is mean , to know how many people come to park on an average basis so that government can get to know about average number of people and maintain the budget of park accordingly.

Part b

We use mean here , i recommend to use mean because it gives more accurate value of average of data when the data is more spreaded like that in our case, where data vary from 1 to 150.

When data is more spreaded , we should use mean which is more accurate.

Mean here is sum of all numbers divided by total numbers.

Mean = (5+3+10+1+2+3+4+3+5+100+4+3+2+4+25+150+3+3+5+4+8+7+10+15+30) ÷ 25

= 409/25= 16.36

Means on an average 16 people come daily

Part c

I'll not recommend median and mode here because the data is wide spreaded and number of observation is not so large, only 25 observations are there and data vary from 1 to 150 that's why median and mode will not give an accurate answer, as median is the number coming in middle of observation when arranged in asending order and mode is the observation which come most of the time.

Median and mode can't give accurate answer.

I prefer mean here.